# To curve, or not to curve: Is curvature-assisted quintessence observationally viable?

### v2

@@ Abstract
Single-field models of accelerated expansion with nearly flat potentials, despite being able to provide observationally viable explanations for the early-time cosmic inflation and the late-time cosmic acceleration, are in strong tension with string theory evidence and the associated de Sitter swampland constraints. It has recently been argued that in an open universe, where the spatial curvature is negative (i.e., with $\Omega_{k}>0$), a new stable fixed point arises, which may lead to viable single-field-based accelerated expansion with an arbitrarily steep potential. Here we show, through a dynamical systems analysis and a Bayesian statistical inference of cosmological parameters, that the additional cosmological solutions based on the new fixed point do not render steep-potential, single-field, accelerated expansion observationally viable. We mainly focus on quintessence models of dark energy, but we also argue that a similar conclusion can be drawn for cosmic inflation.
@@ 1 Introduction
We first provide, in this paper, an overview of the theoretical attempts in the context of string theory, discussing the de Sitter swampland conjecture and its regime of validity. We explain that it is difficult, if not impossible, to obtain models of (early- and late-time) accelerating expansion in the so-called asymptotic regions of moduli space where the theory is under perturbative control. Indeed, in all known examples ever checked one only obtains effective single-field models with exponential scalar potentials that are way too steep to accommodate observationally viable acceleration. This theoretical “evidence” has convinced the majority of string theorists to focus their attention on the bulk of moduli space, beyond the perturbative regime, where computational control is much harder to attain.
@@ 2 String theory and scalar fields
There is a large body of string theory literature devoted to charting the moduli space and the associated potential $V(\vec{\phi})$. The moduli space is often divided into the “bulk” region, where all moduli have a vacuum expectation value (vev) of $\mathcal{O}(1)$, and the “asymptotic” regions, where one or more moduli have a very large vev. Very little is known about the scalar potential in the bulk of moduli space, where there is no control parameter and the theory becomes subject to quantum corrections and non-perturbative effects. In contrast, in asymptotic limits of moduli space, oftentimes we can do a perturbative asymptotic expansion in some of the moduli $\phi^{a}$, and consequently $V(\vec{\phi})$ can be reliably computed. Because of this, there has been a large effort to find de Sitter minima in the asymptotic regions of moduli space. However, despite decades of effort, realizing a de Sitter minimum in these asymptotic regions remains an open question. In fact, every time we compute an asymptotic potential in string theory, it is a sum of exponentials,
@@
which therefore has no minima. This has been checked in many string constructions Grimm et al. (2020); McAllister and Quevedo (2023); Van Riet and Zoccarato (2024), for many moduli, and there are general arguments that this should be the case Castellano et al. (2022, 2023)—there are certainly no counterexamples to (1).
@@
where $d$ is the spacetime dimension, and the norm is taken with respect to the metric in moduli space defined by the moduli kinetic fields, the $\mathsf{G}_{ab}$ in the action
@@
The statement (2) is true in every example ever checked (see, for instance, Refs. Maldacena and Nunez (2001); Hertzberg et al. (2007); Obied et al. (2018); Andriot (2019); Andriot et al. (2020); Calderón-Infante et al. (2023); Shiu et al. (2023a, b); Cremonini et al. (2023); Hebecker et al. (2023); Van Riet (2024); Seo (2024)), and has been conjectured to hold universally Rudelius (2021) in what is known as the “strong de Sitter conjecture” in the “swampland” literature. Weaker versions of the conjecture, where the $2/\sqrt{d-2}$ on the right-hand side is replaced by other values Bedroya and Vafa (2020) or just a generic $\mathcal{O}(1)$ number, have also been put forth, with different general arguments supporting them. We emphasize that all of these statements refer to the asymptotic regions of moduli space, a fact often misunderstood.
In short, from a string theory point of view, it looks like we should leave the easy asymptotic region and start working our way towards the bulk (this is indeed what most proposed de Sitter constructions do; see Refs. Kachru et al. (2003); Danielsson and Dibitetto (2013); Blåbäck et al. (2013); Kachru et al. (2021); Bena et al. (2023); Cicoli et al. (2024) and references therein). However, the recent work Marconnet and Tsimpis (2023); Andriot et al. (2023) revived the interest in asymptotic regions in string theory, showing that one can get acceleration even when $\lambda>\sqrt{2}$, simply by allowing the Universe to have a negative spatial curvature. This fact was already known in the cosmology literature van den Hoogen et al. (1999); Gosenca and Coles (2016), but Refs. Marconnet and Tsimpis (2023); Andriot et al. (2023) emphasized that this would revive the hope that one could obtain accelerated expansion in a fully controlled string theoretic EFT. In the next section, we review this construction in some detail, pointing out a number of features that significantly challenge its cosmological viability.
@@ 3 Curvature to the rescue?
We now describe the dynamical system associated with a universe filled with a single-field quintessence as dark energy, following Refs. van den Hoogen et al. (1999); Gosenca and Coles (2016); Marconnet and Tsimpis (2023); Andriot et al. (2023). We allow the spatial curvature to take a nonzero value, and we also include non-relativistic matter in the analysis with the intention of comparing the system to observational constraints. Although in a fully realistic scenario radiation should also be included, we omit its contribution to the energy budget of the Universe as it becomes negligible long before the accelerated expansion phase begins. We use this approximate dynamical system to draw qualitative conclusions about the viability of quintessence dark energy, before constraining the model in the next section through a full statistical analysis of its parameter space with no approximation.
@@ 3.1 Dynamical system
The action for a canonical scalar field $\phi$ minimally coupled to gravity and in the presence of non-relativistic matter (composed of baryonic and cold dark matter) is given by
@@
where $M_{\rm{Pl}}=1/\sqrt{8\pi G_{\mathrm{N}}}$ is the reduced Plank mass, and $\mathcal{L}_{\rm{m}}$ is the matter Lagrangian. We proceed by assuming the Friedmann-Lemaître-Robertson-Walker (FLRW) metric,
@@
where $H\equiv\dot{a}/a$ is the Hubble expansion rate, and an overdot denotes a derivative with respect to cosmic time $t$. Note that we have assumed a pressureless matter component with the equation of state $w_{\mathrm{m}}=0$. The scalar field $\phi$ follows the equation of motion
@@
For an exponential potential $V(\phi)=V_{0}e^{-\lambda\phi}$, the evolution of the dynamical system can now be described by a system of three coupled differential equations,
@@
Based on the results of Refs. van den Hoogen et al. (1999); Gosenca and Coles (2016), it was emphasized in Refs. Marconnet and Tsimpis (2023); Andriot et al. (2023) that solving these equations could lead to viable cosmological histories with accelerated expansion for $\lambda>\sqrt{2}$ when $k=-1$ (or $\Omega_{k}>0$). This sprung much excitement in the string theory community as these are the kinds of exponential potentials that are easily encountered in string theoretic EFTs. The hope was that this could allow us to embed a realistic cosmology in the asymptotic limit of moduli space, where string theory is under perturbative control. Indeed, in Refs. Marconnet and Tsimpis (2023); Andriot et al. (2023) it was argued that one could obtain an infinite number of $e$-folds of accelerated expansion along some trajectories in phase space, with $\lambda>\sqrt{2}$. Although these statements are true, we now point out a couple of reasons why this scenario might not be as appealing as it appears at first sight.
@@ 3.2 Quality of accelerating phases
Fig. 2 shows the phase space of the dynamical system corresponding to eqs. 18, 19 and 20 for a hyperbolic (or open) universe (i.e., for $k=-1$). The paraboloid in the figure corresponds to the Friedmann equation (14) with $\Omega_{\rm{m}}=0$, above which all possible cosmic trajectories with nonzero $\Omega_{\rm{m}}$ are located. The red and the green dots show the fixed points of the system that can potentially provide trajectories with desired properties, and correspond, respectively, to unstable and attractor solutions. The figure also shows a trajectory for $\lambda=1.6$ (black curve) as an indicative example—the attractor point associated with the trajectory behaves like a node. For a more complete discussion of the fixed points and their stability see Refs. Gosenca and Coles (2016); Marconnet and Tsimpis (2023); Andriot et al. (2023).
Figure 2: Phase space of the dynamical system described by eqs. 18, 19 and 20 for a hyperbolic (or open) universe (i.e., for $k=-1$). The paraboloid corresponds to the Friedmann equation (14) with $\Omega_{\rm{m}}=0$, above which all possible cosmic trajectories with nonzero $\Omega_{\rm{m}}$ are located. Trajectories that pass through the green areas experience accelerating expansion with $w_{\rm{eff}}\in[-1,-0.7]$, while the red regions correspond to $w_{\rm{eff}}=0$ (i.e., matter domination); see eq. 16. The red and green dots show the relevant fixed points of the system and correspond, respectively, to unstable and attractor solutions. For the example of $\lambda=1.6$ the attractor point behaves like a node, as shown by the indicative trajectory (black curve), where the black dot shows the initial conditions.
@@
Since accelerated expansion happens only when $w_{\rm{eff}}<-1/3$, eq. 24 can be rearranged to provide an integral expression for the number of $e$-folds that an accelerated expansion epoch lasts,
@@
approaches 1 exponentially fast; after a few $e$-folds, this cosmology, for all intents and purposes, is indistinguishable from a non-accelerating one. We will show this explicitly in the next section when we constrain the value of $\lambda$ with cosmological observations.
The curvature-assisted acceleration with steep potentials is also in serious trouble when describing an inflationary phase in the early Universe. For single-field, slow-roll cosmic inflation, there is an approximate relation Baumann (2018),
$r\sim 16\,\epsilon\,,$ (27)
@@
Furthermore, the fixed points with $\lambda>\sqrt{2}$ that lead to an arbitrary large number of $e$-folds of accelerated expansion with curvature in Refs. Marconnet and Tsimpis (2023); Andriot et al. (2023) have $z_{*}=-\Omega_{k}$ of $\mathcal{O}(1)$ unless $\lambda$ is very close to $\sqrt{2}$. That means that for large values of $\lambda$, curvature is an $\mathcal{O}(1)$ part of the energy budget of the Universe during the eternally accelerated trajectories. This contrasts with observations, either in inflationary or current era, which pose stringent upper bounds on $\Omega_{k}$.
All of the above show that the eternally accelerated trajectories in Ref. Andriot et al. (2023) with $\lambda>\sqrt{2}$ are of no use for either inflation or to play the role of a quintessence field, since they are ruled out by observations. On the other hand, the current phase of accelerated expansion has only lasted for around $\sim 0.5$ $e$-folds, so one might entertain the possibility that a mild amount of curvature, compatible with observations, may still drive a transient phase of acceleration, even with a steep potential. Determining whether this is really the case can only be done through a careful study of the resulting cosmology, taking into account all observations. The remainder of this paper is devoted to this question.
@@ 4.1 Numerical grid analysis of phase space
Before we confront the curvature-assisted quintessence model with cosmological data, we briefly study the properties of its late-time cosmological solutions through a numerical grid analysis of the dynamical system described by eqs. 18, 19 and 20 and presented in fig. 2. This will further demonstrate that large values of $\lambda$ (i.e., $\lambda>\sqrt{2}$) are incapable of providing observationally viable cosmic histories even with nonzero curvature. Broadly speaking, there are two conditions that any viable solutions must satisfy: 1) provide an extended matter domination phase before the onset of cosmic acceleration, and 2) yield an effective equation of state parameter, $w_{\rm{eff}}$, that is $\sim-0.7$ at the present time, as for the best-fit $\Lambda$CDM model.
@@
Here, we solve the system of differential equations (18)-(20) for different initial conditions, i.e., for different initial values of the three phase space variables $\{x,y,z\}$, and then study the behavior of the cosmological background solution in each case. We do this through a fine grid scan of the entire phase space contained within the paraboloid of fig. 2, defined as $x^{2}+y^{2}-z\leq 1$; see eq. 14. We compute all the trajectories for the different sets of initial conditions and keep only the ones that provide an epoch of $w_{\rm{eff}}\leq-0.7$, i.e., solutions which go through the green regions of fig. 2—we show one example of such solutions for $\lambda=1.6$ in the figure (black curve). Fig. 3 presents the time evolution of $w_{\rm{eff}}$ for all these accelerating solutions. It is clear from the figure that none of the solutions exhibit an epoch of matter domination in the past, i.e., before the accelerating phase of cosmic history. In other words, none of the solutions stay within the red regions of fig. 2 for a sufficiently long time.
@@
Figure 3: Evolution of $w_{\rm{eff}}$ as a function of the number of $e$-folds $N$ for quintessence with $\lambda=1.6$ and for trajectories that provide an epoch of $w_{\rm{eff}}\leq-0.7$, i.e., for those which go through the green regions of fig. 2—these trajectories exhibit late-time accelerated expansion consistent with cosmological observations but do not contain a matter domination epoch.
@@
Figure 4: Upper panel: Evolution of $w_{\phi}$ (equation of state parameter for the scalar field $\phi$), $w_{\rm{eff}}$ (effective equation of state parameter), $\Omega_{\phi}$ (scalar field density parameter), $\Omega_{\rm{m}}$ (non-relativistic matter density parameter), and $\Omega_{k}$ (curvature density parameter) as functions of the number of $e$-folds $N$ for $\lambda=1.6$. The vertical line corresponds to the present time at which $\Omega_{\rm{m}}\sim 0.3$. Lower panel: $w_{\rm{eff}}(N)$ for quintessence with $\lambda=0.5$, $\lambda=1.4$, and $\lambda=1.6$, in comparison with $w_{\rm{eff}}(N)$ for a good-fit $\Lambda$CDM model with $\Omega_{k}$ set to zero. The vertical line corresponds to the present time for the $\Lambda$CDM example.
@@
Having said that, we do find solutions that provide extended matter domination periods, as we have demonstrated in the upper panel of fig. 4 for $\lambda=1.6$ as an example—see how $\Omega_{\rm{m}}$ behaves with time (or number of $e$-folds). However, the problem with these solutions is that none of them provide an effective equation of state parameter $w_{\rm{eff}}$ less than $\sim-0.5$ at the present time (i.e., when $\Omega_{\rm{m}}\sim 0.3$). This means that we do not obtain sufficient acceleration in these cases where matter domination exists. In the lower panel of fig. 4, we present three examples of $w_{\rm{eff}}$ evolution for the curvature-assisted quintessence model, i.e., for $\lambda=\{0.5,1.4,1.6\}$, and compare them with that of a flat $\Lambda$CDM model (i.e., with $\Omega_{k}$ set to zero) that provides a good fit to the data. The figure shows that by increasing $\lambda$ it becomes more and more difficult for the quintessence model to provide a cosmic history that is compatible with cosmological observations, as the current value of $w_{\rm{eff}}$ does not reach the observationally preferred value of $\sim-0.7$, as indicated by our $\Lambda$CDM example.

### v3

@@ Abstract
Single-field models of accelerated expansion with nearly flat potentials, despite being able to provide observationally viable explanations for the early-time cosmic inflation and the late-time cosmic acceleration, are in strong tension with string theory evidence and the associated de Sitter swampland constraints. It has recently been argued that in an open universe, where the spatial curvature is negative (i.e., with $\Omega_{k}>0$), a new stable fixed point arises, which may lead to viable single-field-based accelerated expansion with an arbitrarily steep potential. Here, we show, through a dynamical systems analysis and a Bayesian statistical inference of cosmological parameters, that the additional cosmological solutions based on the new fixed point do not render steep-potential, single-field, accelerated expansion observationally viable. We mainly focus on quintessence models of dark energy, but we also argue that a similar conclusion can be drawn for cosmic inflation.
@@ 1 Introduction
We first provide, in this paper, an overview of the theoretical attempts in the context of string theory, discussing the de Sitter swampland conjecture and its regime of validity. We explain that it is difficult, if not impossible, to obtain models of (early- and late-time) accelerating expansion in the so-called asymptotic regions of moduli space where the theory is under perturbative control. Indeed, in all known examples ever checked, one obtains only effective single-field models with exponential scalar potentials that are way too steep to accommodate observationally viable acceleration. This theoretical “evidence” has convinced the majority of string theorists to focus their attention on the bulk of moduli space, beyond the perturbative regime, where computational control is much harder to attain.
@@ 2 String theory and scalar fields
There is a large body of string theory literature devoted to charting the moduli space and the associated potential $V(\vec{\phi})$. The moduli space is often divided into the “bulk” region, where all moduli have a vacuum expectation value (VEV) of $\mathcal{O}(1)$, and the “asymptotic” regions, where one or more moduli have a very large VEV. Very little is known about the scalar potential in the bulk of moduli space, where there is no control parameter and the theory becomes subject to quantum corrections and nonperturbative effects. In contrast, in asymptotic limits of moduli space, oftentimes we can do a perturbative asymptotic expansion in some of the moduli $\phi^{a}$, and, consequently, $V(\vec{\phi})$ can be reliably computed. Because of this, there has been a large effort to find de Sitter minima in the asymptotic regions of moduli space. However, despite decades of effort, realizing a de Sitter minimum in these asymptotic regions remains an open question. In fact, every time we compute an asymptotic potential in string theory, it is a sum of exponentials:
@@
which, therefore, has no minima. This has been checked in many string constructions Grimm et al. (2020); McAllister and Quevedo (2023); Van Riet and Zoccarato (2024), for many moduli, and there are general arguments that this should be the case Castellano et al. (2022, 2023)—there are certainly no counterexamples to (1).
@@
where $d$ is the spacetime dimension and the norm is taken with respect to the metric in moduli space defined by the moduli kinetic fields, the $\mathsf{G}_{ab}$ in the action
@@
The statement (2) is true in every example ever checked (see, for instance, Refs. Maldacena and Nunez (2001); Hertzberg et al. (2007); Obied et al. (2018); Andriot (2019); Andriot et al. (2020); Calderón-Infante et al. (2023); Shiu et al. (2023a, b); Cremonini et al. (2023); Hebecker et al. (2023); Van Riet (2024); Seo (2024)) and has been conjectured to hold universally Rudelius (2021) in what is known as the “strong de Sitter conjecture” in the “swampland” literature. Weaker versions of the conjecture, where the $2/\sqrt{d-2}$ on the right-hand side is replaced by other values Bedroya and Vafa (2020) or just a generic $\mathcal{O}(1)$ number, have also been put forth, with different general arguments supporting them. We emphasize that all of these statements refer to the asymptotic regions of moduli space, a fact often misunderstood.
In short, from a string theory point of view, it looks like we should leave the easy asymptotic region and start working our way toward the bulk (this is indeed what most proposed de Sitter constructions do; see Refs. Kachru et al. (2003); Danielsson and Dibitetto (2013); Blåbäck et al. (2013); Kachru et al. (2021); Bena et al. (2023); Cicoli et al. (2024) and references therein). However, the recent work Marconnet and Tsimpis (2023); Andriot et al. (2023) revived the interest in asymptotic regions in string theory, showing that one can get acceleration even when $\lambda>\sqrt{2}$, simply by allowing the Universe to have a negative spatial curvature. This fact was already known in the cosmology literature van den Hoogen et al. (1999); Gosenca and Coles (2016), but Refs. Marconnet and Tsimpis (2023); Andriot et al. (2023) emphasized that this would revive the hope that one could obtain accelerated expansion in a fully controlled string theoretic EFT. In the next section, we review this construction in some detail, pointing out a number of features that significantly challenge its cosmological viability.
@@ 3 Curvature to the rescue?
We now describe the dynamical system associated with a universe filled with a single-field quintessence as dark energy, following Refs. van den Hoogen et al. (1999); Gosenca and Coles (2016); Marconnet and Tsimpis (2023); Andriot et al. (2023). We allow the spatial curvature to take a nonzero value, and we also include nonrelativistic matter in the analysis with the intention of comparing the system to observational constraints. Although in a fully realistic scenario radiation should also be included, we omit its contribution to the energy budget of the Universe as it becomes negligible long before the accelerated expansion phase begins. We use this approximate dynamical system to draw qualitative conclusions about the viability of quintessence dark energy, before constraining the model in the next section through a full statistical analysis of its parameter space with no approximation.
@@ 3.1 Dynamical system
The action for a canonical scalar field $\phi$ minimally coupled to gravity and in the presence of nonrelativistic matter (composed of baryonic and cold dark matter) is given by
@@
where $M_{\rm{Pl}}=1/\sqrt{8\pi G_{\mathrm{N}}}$ is the reduced Plank mass, and $\mathcal{L}_{\rm{m}}$ is the matter Lagrangian. We proceed by assuming the Friedmann-Lemaître-Robertson-Walker metric
@@
where $H\equiv\dot{a}/a$ is the Hubble expansion rate and an overdot denotes a derivative with respect to cosmic time $t$. Note that we have assumed a pressureless matter component with the equation of state $w_{\mathrm{m}}=0$. The scalar field $\phi$ follows the equation of motion
@@
For an exponential potential $V(\phi)=V_{0}e^{-\lambda\phi}$, the evolution of the dynamical system can now be described by a system of three coupled differential equations:
@@
Based on the results of Refs. van den Hoogen et al. (1999); Gosenca and Coles (2016), it was emphasized in Refs. Marconnet and Tsimpis (2023); Andriot et al. (2023) that solving these equations could lead to viable cosmological histories with accelerated expansion for $\lambda>\sqrt{2}$ when $k=-1$ (or $\Omega_{k}>0$). This sprung much excitement in the string theory community, as these are the kinds of exponential potentials that are easily encountered in string theoretic EFTs. The hope was that this could allow us to embed a realistic cosmology in the asymptotic limit of moduli space, where string theory is under perturbative control. Indeed, in Refs. Marconnet and Tsimpis (2023); Andriot et al. (2023), it was argued that one could obtain an infinite number of $e$-folds of accelerated expansion along some trajectories in phase space, with $\lambda>\sqrt{2}$. Although these statements are true, we now point out a couple of reasons why this scenario might not be as appealing as it appears at first sight.
@@ 3.2 Quality of accelerating phases
Figure 2 shows the phase space of the dynamical system corresponding to Eqs. 18, 19 and 20 for a hyperbolic (or open) universe (i.e., for $k=-1$). The paraboloid in the figure corresponds to the Friedmann equation (14) with $\Omega_{\rm{m}}=0$, above which all possible cosmic trajectories with nonzero $\Omega_{\rm{m}}$ are located. The red and the green dots show the fixed points of the system that can potentially provide trajectories with desired properties and correspond, respectively, to unstable and attractor solutions. The figure also shows a trajectory for $\lambda=1.6$ (black curve) as an indicative example—the attractor point associated with the trajectory behaves like a node. For a more complete discussion of the fixed points and their stability see Refs. Gosenca and Coles (2016); Marconnet and Tsimpis (2023); Andriot et al. (2023).
Figure 2: Phase space of the dynamical system described by Eqs. 18, 19 and 20 for a hyperbolic (or open) universe (i.e., for $k=-1$). The paraboloid corresponds to the Friedmann equation (14) with $\Omega_{\rm{m}}=0$, above which all possible cosmic trajectories with nonzero $\Omega_{\rm{m}}$ are located. Trajectories that pass through the green areas experience accelerating expansion with $w_{\rm{eff}}\in[-1,-0.7]$, while the red regions correspond to $w_{\rm{eff}}=0$ (i.e., matter domination); see Eq. 16. The red and green dots show the relevant fixed points of the system and correspond, respectively, to unstable and attractor solutions. For the example of $\lambda=1.6$, the attractor point behaves like a node, as shown by the indicative trajectory (black curve), where the black dot shows the initial conditions.
@@
Since accelerated expansion happens only when $w_{\rm{eff}}<-1/3$, Eq. 24 can be rearranged to provide an integral expression for the number of $e$-folds that an accelerated expansion epoch lasts:
@@
approaches 1 exponentially fast; after a few $e$-folds, this cosmology, for all intents and purposes, is indistinguishable from a nonaccelerating one. We will show this explicitly in the next section when we constrain the value of $\lambda$ with cosmological observations.
The curvature-assisted acceleration with steep potentials is also in serious trouble when describing an inflationary phase in the early Universe. For single-field, slow-roll cosmic inflation, there is an approximate relation Baumann (2018)
$r\sim 16\,\epsilon$ (27)
@@
Furthermore, the fixed points with $\lambda>\sqrt{2}$ that lead to an arbitrary large number of $e$-folds of accelerated expansion with curvature in Refs. Marconnet and Tsimpis (2023); Andriot et al. (2023) have $z_{*}=-\Omega_{k}$ of $\mathcal{O}(1)$ unless $\lambda$ is very close to $\sqrt{2}$. That means that for large values of $\lambda$, curvature is an $\mathcal{O}(1)$ part of the energy budget of the Universe during the eternally accelerated trajectories. This contrasts with observations, in either the inflationary or current era, which pose stringent upper bounds on $\Omega_{k}$.
All of the above show that the eternally accelerated trajectories in Ref. Andriot et al. (2023) with $\lambda>\sqrt{2}$ are of no use for either inflation or to play the role of a quintessence field, since they are ruled out by observations. On the other hand, the current phase of accelerated expansion has lasted for only around $\sim 0.5$ $e$-folds, so one might entertain the possibility that a mild amount of curvature, compatible with observations, may still drive a transient phase of acceleration, even with a steep potential. Determining whether this is really the case can be done only through a careful study of the resulting cosmology, taking into account all observations. The remainder of this paper is devoted to this question.
@@ 4.1 Numerical grid analysis of phase space
Before we confront the curvature-assisted quintessence model with cosmological data, we briefly study the properties of its late-time cosmological solutions through a numerical grid analysis of the dynamical system described by Eqs. 18, 19 and 20 and presented in Fig. 2. This will further demonstrate that large values of $\lambda$ (i.e., $\lambda>\sqrt{2}$) are incapable of providing observationally viable cosmic histories even with nonzero curvature. Broadly speaking, there are two conditions that any viable solutions must satisfy: (i) provide an extended matter domination phase before the onset of cosmic acceleration, and (ii) yield an effective equation of state parameter $w_{\rm{eff}}$, that is $\sim-0.7$ at the present time, as for the best-fit $\Lambda$CDM model.
@@
Here, we solve the system of differential equations (18)-(20) for different initial conditions, i.e., for different initial values of the three phase space variables $\{x,y,z\}$, and then study the behavior of the cosmological background solution in each case. We do this through a fine grid scan of the entire phase space contained within the paraboloid in Fig. 2, defined as $x^{2}+y^{2}-z\leq 1$; see Eq. 14. We compute all the trajectories for the different sets of initial conditions and keep only the ones that provide an epoch of $w_{\rm{eff}}\leq-0.7$, i.e., solutions which go through the green regions in Fig. 2—we show one example of such solutions for $\lambda=1.6$ in the figure (black curve). Figure 3 presents the time evolution of $w_{\rm{eff}}$ for all these accelerating solutions. It is clear from the figure that none of the solutions exhibit an epoch of matter domination in the past, i.e., before the accelerating phase of cosmic history. In other words, none of the solutions stay within the red regions in Fig. 2 for a sufficiently long time.
@@
Figure 3: Evolution of $w_{\rm{eff}}$ as a function of the number of $e$-folds $N$ for quintessence with $\lambda=1.6$ and for trajectories that provide an epoch of $w_{\rm{eff}}\leq-0.7$, i.e., for those which go through the green regions in Fig. 2—these trajectories exhibit late-time accelerated expansion consistent with cosmological observations but do not contain a matter domination epoch.
@@
Figure 4: Upper panel: evolution of $w_{\phi}$ (equation of state parameter for the scalar field $\phi$), $w_{\rm{eff}}$ (effective equation of state parameter), $\Omega_{\phi}$ (scalar field density parameter), $\Omega_{\rm{m}}$ (nonrelativistic matter density parameter), and $\Omega_{k}$ (curvature density parameter) as functions of the number of $e$-folds $N$ for $\lambda=1.6$. The vertical line corresponds to the present time at which $\Omega_{\rm{m}}\sim 0.3$. Lower panel: $w_{\rm{eff}}(N)$ for quintessence with $\lambda=0.5$, $\lambda=1.4$, and $\lambda=1.6$, in comparison with $w_{\rm{eff}}(N)$ for a good-fit $\Lambda$CDM model with $\Omega_{k}$ set to zero. The vertical line corresponds to the present time for the $\Lambda$CDM example.
@@
Having said that, we do find solutions that provide extended matter domination periods, as we have demonstrated in the upper panel in Fig. 4 for $\lambda=1.6$ as an example—see how $\Omega_{\rm{m}}$ behaves with time (or number of $e$-folds). However, the problem with these solutions is that none of them provide an effective equation of state parameter $w_{\rm{eff}}$ less than $\sim-0.5$ at the present time (i.e., when $\Omega_{\rm{m}}\sim 0.3$). This means that we do not obtain sufficient acceleration in these cases where matter domination exists. In the lower panel in Fig. 4, we present three examples of $w_{\rm{eff}}$ evolution for the curvature-assisted quintessence model, i.e., for $\lambda=\{0.5,1.4,1.6\}$, and compare them with that of a flat $\Lambda$CDM model (i.e., with $\Omega_{k}$ set to zero) that provides a good fit to the data. The figure shows that by increasing $\lambda$ it becomes more and more difficult for the quintessence model to provide a cosmic history that is compatible with cosmological observations, as the current value of $w_{\rm{eff}}$ does not reach the observationally preferred value of $\sim-0.7$, as indicated by our $\Lambda$CDM example.